CRY+VIS Interface Meeting
Date: 2022/1/21 13:00-13:50
Participants: Takahashi, Hirata, Yokozawa, Aso, Ikeda, Kimura, Tomaru, Miyoki, Ushiba, Sato, Akutsu
- IX: Health check is ongoing. Glitch in the F3 signals is still under investigation. Windshield for the MN oplev should be modified.
- EX: Health check is ongoing. Leake check was scheduled on 1/25.
- IY: Accelerometers were installed.
- EY: IP H2 actuator and Photosensor V3 were repaired. Preparing for the accelerometers. Oplev setting was scheduled on 1/26-27.
- Replacement of flip cables: SRM was finished. Cables for Type-C will be delivered in the end of Jan.
- TM height setting: ETMY, ITMY.
- Oplev: Setup of the TM Oplev in IY after the TM alignment in late Feb.
- ACC: Preparation for EY.
- Commissioning in air: Return of GRY. Alignment of SR3/ETMY.
- Commissioning in vacuum: Go to the stable lock.
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: EXV (transportation on 2/7, implementation on 2/8-11)
- Heater control: Test in IXV/EXV is planed. Repaired power controller was delivered.
- Power source for Oplev: Next is IY
- Tunnel for the IYC booth: Scheduled on 1/31.
- TMS: Scheduled on 2/14-16.
Date: 2022/1/21 14:00-14:20
Participants: Takahashi, Ushiba, Fabian, Terrence, Ishizaki, Tanaka, Aso
- Installation of the accelerometer in IY.
- Repair of the IP H2 actuator and the cable for the V3 photosensor in EY.
- Preparation for the accelerometers in EY.
- [Terrence] Tower control was implemented.
- Guardian was made to go to ALIGNED state with MN OPTICALIGN offsets.
- MN to TM DC control was implemented (Not posted yet).
- I need to modify Guardian to engage this, but one of the photosensor is not working properly so there's no damping at the MN stage so I postponed this.
- We finished replacing the flip adapters in SRM
- I compiled the results of the cable health check in one spreadsheet in ordert to compare values that should be similar, and I found two possible problems:
- In SR2, the resistance of IM-V3 coil was higher than expected by about 3 Ohm: we checked and realized the additional resistance comes from a extension cable, but the cable is healthy.
- In the BS, the a resistance of the F0 FR stepper motor is higher by several Ohm, we need to check next week.
- In the BS F1 LVDT, we changed the polrity of the LVDT but we didn't confirmed it worked until this week: positive actuation, which moves the keystone upwards, produces a positive LVDT output.
- We tried to calibrate the BS F1 LVDT today, but we were not able to align the displacement sensor properly. The holder can only be adjusted in one angular DoF, but we need to be able to adjust in teo angular DoF.
I compiled the results of the cable health check and found some potential problems, which we aim to fix next week.
- PR2 BF FR, pins 1-6, P1-7: the resistance is 9.7, which is higher than the values in other BF FRs in PR suspensions (8.8 Ohm maximum).
- PR3 BF V3, pins 1-6, P4-7: the resistance is 13.7 Ohm, which is slightly higher than 11.6 Ohm measured in other LVDTs (approximately).
- PR3 Traverser: there are several resistance values that are several Ohm higher than expected:
- R1, P-2-2: pins 1-2, 1-3, 2-3, 2-6, 3-6, 3-7, 6-7.
- R2, P4-2: pins 1-6, 2-6, 3-6, 6-7.
- Hirata-san and Nakagaki-san removed some unused cables that were hanging close to PR2.
- Hirata-san designed cable clamps to be installed at the upper part Type-Bp suspensions.
- Type-A team: Installation of the accelerometers in EY, Height adjustment of ETMY.
- Type-B team: PR cables, BS SF calibration.
(Date in Kamioka)
The week from 1/24
- Takahashi: 1/25-28
- Sato: 1/25pm-28 or 1/25-27
- Hirata: 1/25-28
The week from 1/31
- Takahashi: 1/31-2/4
- Sato: 1/31-2/4
- Hirata: 1/31-2/4